# TokaMark: A Comprehensive Benchmark for MAST Tokamak Plasma Models

### v1

@@ 4.1. Model Description
Encoders. For each input variable $v_{i}\in(L_{t}^{(0,v_{i})}\times L_{h}^{(0,v_{i})}\times L_{w}^{(0,v_{i})})$, an independent convolutional encoder is instantiated according to its modality: 1D convolutions for time series, 2D convolutions for profiles, and 3D convolutions for videos signals. Each encoder consists of a stack of $N$ convolutional layers with kernel size $K$, stride $s$, and padding $p$, followed by ReLU activations, max-pooling, and batch normalization. The feature maps are then flattened into a latent vector. The size of the feature map $L_{d}^{(\ell,v_{i})}$ along dimension $d\in(t,h,w)$ after encoder layer $\ell\in[1,N]$ and the flattened latent representation size $L^{(v_{i})}$ for variable $v_{i}$ can be computed as follows:
(5) $L_{d}^{(\ell,v_{i})}$ $=\left\lfloor\frac{1}{2}\left\lfloor\frac{L_{d}^{(\ell-1,v_{i})}+2p-K}{s}\right\rfloor+p-\frac{1}{2}\right\rfloor+1$
(6) $L^{(v_{i})}$ $=2^{N-1}DL_{t}^{(N,v_{i})}L_{h}^{(N,v_{i})}L_{w}^{(N,v_{i})}$
Figure 2. Multi-branch convolutional encoder–decoder
Latent fusion backbone. The latent embeddings produced by all encoder branches are concatenated and passed through a shared backbone. This backbone consists of a sequence of linear layers each followed by ReLU activation and dropout, progressively reducing the dimensionality back to $D$. This shared representation provides a compact summary of all input data and connects the encoders and decoders blocks.
@@
(7) $L_{d}^{(\ell,v_{o})}$ $=\left\lceil s\cdot(L_{d}^{(\ell+1,v_{o})}-1)-2p+K+o_{p}\right\rceil$
(8) $L^{(v_{o})}$ $=2^{N-1}DL_{t}^{(0,v_{o})}L_{h}^{(0,v_{o})}L_{w}^{(0,v_{o})}$
@@ 4.2. Experimental Settings
Parameter Settings. For our architecture, we adopt a latent embedding dimension of $D=16$, with $N=3$ convolutional layers per encoder and decoder blocks. Each convolution uses a kernel size $K=3$, stride $s=3$, padding $p=1$ and output padding $o_{p}=1$. The same architectural hyperparameters — number of layers, kernel size, stride, and padding — are applied consistently across all tasks, establishing a unified baseline for the benchmark.
Data Preprocessing. We run a set of model specific preprocessing steps: first, we replace NaN values with zeros, second, we standardize data using signal-computed zero mean and unit variance scaler. This ensures numerical stability and helps to improve convergence during training. The training and validation data is down-sampled by taking window samples with a stride of $0.005$ seconds for Markovian tasks and $0.025$ seconds for non-Markovian tasks. Furthermore, because our architecture requires a fixed-length context, all input windows are truncated to a duration of 50 ms for non-Markovian tasks.
Training Procedure. We train our model on mini-batches containing data from 32 shots using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and early stopping with a patience of 10 epochs which terminates training if the validation loss does not improve over this period to prevent overfitting. We also employ a multi-output mean squared error loss, which averages the loss across all outputs of the model. Table 3 summarizes the size of models per task, computed from the encoder and decoder feature maps.
@@ 4.3. Experimental Results
The results of the baseline models evaluation on TokaMark tasks are presented in Table 3 containing task level scores and group averages. Those results demonstrate a clear distinction between tasks complexities: the model performs well for equilibrium reconstruction (Group 1) and magnetics dynamics (Group 2) tasks with group-level scores below 0.17 but degrades for profile dynamics (Group 3) and MHD activites (Group 4) tasks confirming that, as expected, these are more difficult tasks. Notably, even within the same group tasks scores vary substantially. For instance, plasma boundary reconstruction and forecasting tasks 1-2 and 2-2 are resolved much better then their counterparts within the respective groups. On the other side, the error score for task 4-5 is by far the worst and exceeds unity, suggesting the corresponding signals are poorly constrained or inadequately represented.
These results should be interpreted with a consideration to the nature of a baseline model: the architecture is generic, without physics-informed priors or task-specific tuning. The scores reported – especially those for profile dynamics and MHD activities tasks – highlight intrinsic benchmark difficulty rather than flaws in optimization. This baseline establishes a realistic lower bound and identifies areas for improvement, such as physics-aware modeling, task-adaptive weighting, or specialized architectures that should be considered by the users of TokaMark.
Table 3. TaskScore and GroupScore for TokaMark.
Task TaskScore Model size (M)

### v2

@@ 4.1. Model Description
Encoders. For each input variable $v_{i}\in(L_{t}^{(0,v_{i})}\times L_{h}^{(0,v_{i})}\times L_{w}^{(0,v_{i})})$, an independent convolutional encoder is instantiated according to its modality: 1D convolutions for time series, 2D convolutions for profiles, and 3D convolutions for video signals. Each encoder consists of a stack of $N$ convolutional layers with kernel size $K$, stride $s$, and padding $p$, followed by ReLU activations, max-pooling, and batch normalization. The feature maps are then flattened into a latent vector. The size of the feature map $L_{d}^{(\ell,v_{i})}$ along dimension $d\in(t,h,w)$ after encoder layer $\ell\in[1,N]$ and the flattened latent representation size $L^{(v_{i})}$ for variable $v_{i}$ can be computed as follows:
(5) $L_{d}^{(\ell,v_{i})}$ $=\left\lfloor\frac{1}{2}\left\lfloor\frac{L_{d}^{(\ell-1,v_{i})}+2p-K}{s}\right\rfloor+p-\frac{1}{2}\right\rfloor+1,$
(6) $L^{(v_{i})}$ $=2^{N-1}DL_{t}^{(N,v_{i})}L_{h}^{(N,v_{i})}L_{w}^{(N,v_{i})}.$
Figure 2. Multi-branch convolutional encoder–decoder.
Latent fusion backbone. The latent embeddings produced by all encoder branches are concatenated and passed through a shared backbone. This backbone consists of a sequence of linear layers, each followed by ReLU activation and dropout, progressively reducing the dimensionality back to $D$. This shared representation provides a compact summary of all input data and connects the encoder and decoder blocks.
@@
(7) $L_{d}^{(\ell,v_{o})}$ $=\left\lceil s\cdot(L_{d}^{(\ell+1,v_{o})}-1)-2p+K+o_{p}\right\rceil,$
(8) $L^{(v_{o})}$ $=2^{N-1}DL_{t}^{(0,v_{o})}L_{h}^{(0,v_{o})}L_{w}^{(0,v_{o})}.$
@@ 4.2. Experimental Settings
Parameter Settings. For our architecture, we adopt a latent embedding dimension of $D=16$, with $N=3$ convolutional layers per encoder and decoder blocks. Each convolution uses a kernel size $K=3$, stride $s=3$, padding $p=1$, and output padding $o_{p}=1$. The same architectural hyperparameters—number of layers, kernel size, stride, and padding—are applied consistently across all tasks.
Data Preprocessing. We run a set of model-specific preprocessing: first, we replace NaN values with zeros, and then we standardize data using signal-computed zero mean and unit variance scaler. This ensures numerical stability and helps to improve training convergence. The training and validation data is downsampled by taking samples with a stride of $0.005$ seconds for Markovian tasks, and $0.025$ seconds for non-Markovian tasks. Furthermore, because our architecture requires a fixed-length context, inputs for non-Markovian tasks are truncated to a duration of $50$ ms.
Training Procedure. We train our model on mini-batches containing data from 32 shots using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and early stopping with a patience of 10 epochs. We also employ a multi-output mean squared error loss, which averages the loss across all outputs of the model. Table 3 summarizes models sizes per each task.
@@ 4.3. Experimental Results
The results of the baseline model evaluation on TokaMark tasks are presented in Table 3. Those results demonstrate a clear distinction between tasks complexities: the model performs well for equilibrium reconstruction (Group 1) and magnetics dynamics (Group 2) tasks with the group-level scores below 0.17, but degrades for profile dynamics (Group 3) and MHD activities (Group 4) tasks, confirming these are more difficult tasks. Notably, even within the same group, tasks scores vary substantially. For instance, plasma boundary reconstruction and forecasting tasks Task 1-2 and Task 2-2 are resolved much better than their counterparts within the respective groups. On the other side, the error score for Task 4-5 is by far the worst and exceeds unity, suggesting the corresponding signals are poorly constrained or inadequately represented. Finally, we provide signal level error metrics in the appendix in Table A.2.
These results should be interpreted with a consideration to the nature of a baseline model: the architecture is generic, without physics-informed priors or task-specific tuning. The scores reported—especially those for profile dynamics and MHD activity tasks—highlight intrinsic benchmark difficulty rather than flaws in optimization. Nevertheless, this baseline establishes a realistic lower bound and identifies areas for improvement.
Table 3. Task and Group errors for TokaMark  tasks.
$\textbf{Task/Group}_{\mathrm{\textbf{NRMSE}}}$ Parameters(M)
@@ 5. Conclusions
Appendix A Full Set of Signals
Table A.1 presents the complete set of signals used in TokaMark. The table is aligned with the taxonomy in Table 1 and specifies signals within each category and subcategory. We note, that both the data used by TokaMark as well as the FAIR-MAST data are stored in per-shot zarr files. Within each zarr file, signals are organized into groups (which are similar, though not identical, to the categories listed in Table A.1). We therefore provide each signal’s full identifier in the format ”group name”-”signal name”.
Appendix B Signal Level Errors
In addition to the task- and group-level errors we also report more granular signal-level errors in Table A.2. Moreover, that table includes not only NRMSE-based signal errors but also RMSE-, MAE- and NMAE-based errors, all computed analogously.
Table A.1. Full set of signals used in TokaMark.
Category Subcategory/Signals Full signal name
Magnetics Flux loops magnetics-flux_loop_flux
Pickup coils magnetics-[b_field_pol_probe_ccbv_field, b_field_pol_probe_obr_field, b_field_pol_probe_obv_field]
Saddle coils magnetics-b_field_tor_probe_saddle_voltage
Kinetics Thomson scattering thomson_scattering-[t_e, n_e]
Interferometer interferometer-n_e_line
[…]
soft X-ray soft_x_rays-[horizontal_cam_lower, horizontal_cam_upper]
Fast magnetics Mirnov coils magnetics-[b_field_tor_probe_cc_field, b_field_pol_probe_omv_voltage]
Currents Poloidal field coil currents pf_active-[coil_current, solenoid_current]
Plasma current summary-ip
Voltages Poloidal field coil voltages pf_active-coil_voltage
References Reference plasma current pulse_schedule-i_plasma
Reference plasma density pulse_schedule-n_e_line
Fueling NBI power summary-power_nbi
Gas puffing gas_injection-total_injected
Equilibrium Shape parameters equilibrium-[elongation, elongation_axis, triangularity_upper, triangularity_lower, x_point_r, x_point_z, minor_radius, magnetic_axis_r, magnetic_axis_z]
$J_{\text{tor}}$ metrics equilibrium-[q95, beta_tor, beta_pol, beta_normal, bvac_rmag, bphi_rmag]
plasma boundary equilibrium-[lcfs_r, lcfs_z]
flux map equilibrium-psi
Table A.2. Signal errors across all tasks and groups.
Task Full signal name $\textbf{Signal}_{\mathrm{\textbf{RMSE}}}$ $\textbf{Signal}_{\mathrm{\textbf{MAE}}}$ $\textbf{Signal}_{\mathrm{\textbf{NRMSE}}}$ $\textbf{Signal}_{\mathrm{\textbf{NMAE}}}$
Group 1 1-1 equilibrium-beta_normal 0.3686 0.2664 0.2734 0.1976
equilibrium-beta_pol 0.0870 0.0609 0.2668 0.1868
equilibrium-beta_tor 0.8385 0.6512 0.2415 0.1876
equilibrium-bphi_rmag 0.0293 0.0251 0.1497 0.1281
equilibrium-bvac_rmag 0.0230 0.0213 0.1367 0.1270
equilibrium-elongation 0.0501 0.0414 0.2342 0.1936
equilibrium-elongation_axis 0.0564 0.0472 0.2676 0.2240
equilibrium-magnetic_axis_r 0.0147 0.0117 0.1557 0.1237
equilibrium-magnetic_axis_z 0.0105 0.0086 0.1429 0.1173
equilibrium-minor_radius 0.0114 0.0091 0.1805 0.1437
equilibrium-q95 0.5950 0.4729 0.1749 0.1390
equilibrium-triangularity_lower 0.0223 0.0181 0.2277 0.1851
equilibrium-triangularity_upper 0.0220 0.0178 0.2155 0.1744
equilibrium-x_point_r 0.1099 0.0799 0.1431 0.1041
equilibrium-x_point_z 0.0926 0.0599 0.0130 0.0084
1-2 equilibrium-lcfs_r 0.0211 0.0147 0.0559 0.0390
equilibrium-lcfs_z 0.0262 0.0197 0.0404 0.0303
1-3 equilibrium-psi 0.0142 0.0050 0.2552 0.0901
Group 2 2-1 equilibrium-beta_normal 0.3102 0.2210 0.2301 0.1639
equilibrium-beta_pol 0.0734 0.0509 0.2252 0.1562
equilibrium-beta_tor 0.7493 0.5718 0.2158 0.1647
equilibrium-bphi_rmag 0.0225 0.0178 0.1152 0.0911
equilibrium-bvac_rmag 0.0181 0.0163 0.1079 0.0969
equilibrium-elongation 0.0460 0.0373 0.2147 0.1742
equilibrium-elongation_axis 0.0533 0.0441 0.2527 0.2090
equilibrium-magnetic_axis_r 0.0138 0.0109 0.1462 0.1155
equilibrium-magnetic_axis_z 0.0102 0.0079 0.1393 0.1080
equilibrium-minor_radius 0.0106 0.0084 0.1677 0.1324
equilibrium-q95 0.5151 0.4151 0.1514 0.1220
equilibrium-triangularity_lower 0.0214 0.0170 0.2187 0.1734
equilibrium-triangularity_upper 0.0214 0.0168 0.2093 0.1641
equilibrium-x_point_r 0.0976 0.0656 0.1272 0.0854
equilibrium-x_point_z 0.0870 0.0557 0.0122 0.0078
pf_active-coil_current 535.13 374.30 0.1057 0.0739
pf_active-solenoid_current 913.93 749.73 0.0726 0.0595
summary-ip 26035.74 21395.01 0.0803 0.0660
2-2 equilibrium-lcfs_r 0.0212 0.0149 0.0561 0.0393
equilibrium-lcfs_z 0.0307 0.0230 0.0473 0.0355
2-3 equilibrium-psi 0.0096 0.0037 0.1724 0.0663
Group 3 3-1 thomson_scattering-n_e $4.73\times 10^{18}$ $3.43\times 10^{18}$ 0.3307 0.2399
thomson_scattering-t_e 107.00 71.96 0.3196 0.2149
3-2 soft_x_rays-horizontal_cam_lower 0.00719 0.00343 0.2382 0.1136
soft_x_rays-horizontal_cam_upper 0.00735 0.00355 0.2717 0.1313
spectrometer_visible-filter_spectrometer_dalpha_voltage 0.2172 0.1226 0.4830 0.2726
3-3 equilibrium-beta_normal 0.4389 0.2981 0.3256 0.2211
equilibrium-beta_pol 0.0987 0.0628 0.3028 0.1926
equilibrium-beta_tor 0.9948 0.7296 0.2865 0.2101
thomson_scattering-n_e $7.35\times 10^{18}$ $5.69\times 10^{18}$ 0.5143 0.3984
thomson_scattering-t_e 125.35 94.33 0.3744 0.2817
Group 4 4-1 soft_x_rays-horizontal_cam_lower 0.00962 0.00523 0.3184 0.1731
soft_x_rays-horizontal_cam_upper 0.01003 0.00562 0.3705 0.2077
4-2 soft_x_rays-horizontal_cam_lower 0.00971 0.00488 0.3215 0.1616
soft_x_rays-horizontal_cam_upper 0.00922 0.00485 0.3406 0.1792
4-3 equilibrium-magnetic_axis_z 0.0198 0.0171 0.2702 0.2341
4-4 summary-ip 139237.10 91351.51 0.4292 0.2816
4-5 magnetics-b_field_pol_probe_omv_voltage 0.1595 0.0683 0.9371 0.4014
magnetics-b_field_tor_probe_cc_field $2.15\times 10^{-5}$ $5.95\times 10^{-6}$ 1.0734 0.2972